Two pendulums of time periods 3s and 7 s respectively start oscillating simultaneously from two opposite extreme positions. After how much time they will be in the same phase?

Grade:12

1 Answers

Joy Ferreira
5 years ago
y1=Asin(w1t+π)  {π because both particles are at extreme position}                                                                                                 y2= Asinw2t                                                                                        since they are in same phase, y1=y2                                             Asin(w1t+π)=Asinw2t                                                                                                                w1t +π=w2t                                               π=w2t-w1t  =  2πt/T1 -2πt/T2.                                                                      π=2πt[1/T1 - 1/T2] .                                                                          1=2[1/3-1/7]t.  =   (2×4/21) t .                                                     1=(8/21) t                                                                                             t=21/8